OpenStudy (anonymous):

If 5.85 grams of cobalt metal react with 15.8 grams of silver nitrate, how many grams of silver metal can be formed and how many grams of the excess reactant will be left over when the reaction is complete? Show all of your work. unbalanced equation: Co + AgNO3 “yields”/ Co(NO3)2 + Ag I balanced it Co + 2AgNO3 --> Co(NO3)2 + 2Ag

OpenStudy (somy):

Step 1 find moles of each

OpenStudy (somy):

you have masses given for you

OpenStudy (somy):

mole= mass/ molar mass

OpenStudy (anonymous):

Moles of Co and AgNO3? I know the math by the way i just need to be sure

OpenStudy (anonymous):

?

OpenStudy (somy):

yes

OpenStudy (somy):

do the calculation here and show what you are getting so that we confirm it :) and then move to next step :D

OpenStudy (anonymous):

Ok... Co = 5.85 g / 58.933 g/mol= 0.0993 AgNO3 = 15.8 g / 169.8717 g/mol= 0.0930? I think i did this one wrong

OpenStudy (anonymous):

AgNO3 is the LR

OpenStudy (anonymous):

Hello?

OpenStudy (somy):

sorry yes im here why you think you did that one wrong?

OpenStudy (anonymous):

Never mind. i checked it aging

OpenStudy (somy):

alright well i think you did it right

OpenStudy (anonymous):

K what's next?

OpenStudy (somy):

Which one is limiting now?

OpenStudy (anonymous):

AgNO3

OpenStudy (somy):

alright good now Limiting reagent as you know finishes first

OpenStudy (anonymous):

k

OpenStudy (somy):

that is why reaction/amount of product depends on limiting reagent

OpenStudy (anonymous):

Wait am i supposed to do something else now? sorry i just have not been doing this for a while and need a refresh

OpenStudy (somy):

thats why we use mole of limiting reagent and we look at the ratio (meaning coefficients)

OpenStudy (anonymous):

right

OpenStudy (abmon98):

Number of Moles=Mass(g)/Molar Mass(g/mol) Co + 2AgNO3-->Co(NO3)2 + 2 Ag 5.85/58.9=0.0993 moles of Cobalt 15.8/169.8=0.093 moles of Silver Nitrate Look at the mole ratio of Co to AgNO3 1:2 0.0993:x moles AgNO3 required = 0.0993 x 2 = 0.199 limiting reagent is the reactant that is completely used up in a reaction and thus determines when the reaction stops. The number of moles required(calculated) is greater than the present number of mole therefore AgNO3 moles is the limiting reagent. How many grams of silver metal can be formed? Ratio of moles of 2AgNO3 to Ag 2:2 they share the same number of moles 0.093*107.8= How many grams of the excess reactant will be left over when the reaction is complete? Look at the mole Ratio of Co to AgNO3 1:2 x :0.0930 x=0.0930*1/2=0.0465 0.0993-0.0465=Mass of left over reactant

OpenStudy (somy):

Awesome job as always!!! @Abmon98 sorry my pc internet went off <.<

OpenStudy (abmon98):

Thank you :D No Problem